# LagCeresa.bib

### v1

@@ Acknowledgements
The notion of Lagrangian cobordism was originally introduced by Arnold in [Arnold_Lag&Leg_cobordisms].
@@
Two finite sets of Lagrangians $(L_{1},\dots,L_{k^{+}})$ and $(L^{\prime}_{1},\dots,L^{\prime}_{k^{-}})$ in $X$ are Lagrangian cobordant if there exists a Lagrangian $V$ in $X\times\mathbb{C}$ and a compact subset $K\subset X\times\mathbb{C}$ such that
@@
for pairwise distinct real numbers $\{x_{i}\}$ and $\{x^{\prime}_{j}\}$. The Lagrangian cobordism group of $X$, $\mathcal{C}ob(X)$, is then
$\mathcal{C}ob(X):=\bigoplus_{L\in X}\mathbb{Z}\cdot L/\sim,$
@@
This definition can be modified to fit the desired set-up by adding assumptions on the finite sets of Lagrangians and the cobordisms themselves with a compatibility condition on the ends; in this way one can upgrade Lagrangians to Lagrangian branes, require exactness, monotonicity, etc. Most relevant to our considerations is the oriented Lagrangian cobordism group, which we denote by $\mathcal{C}ob^{or}(X)$.
A slight variation of Lagrangian cobordisms are cylindrical Lagrangian cobordisms, introduced by Sheridan–Smith in [nick&ivan1]. These differ from Definition 2.1 in that $V$ is a Lagrangian in $X\times\mathbb{C}^{*}$, which outside of a compact subset in $\mathbb{C}^{*}$ carries trivially fibered finite sets of Lagrangians in $X$ over radial rays. This variation appears naturally from mirror symmetry, where cylindrical cobordisms provide the more natural mirror to rational equivalence.
@@
Let $L\subset X$ be a Lagrangian, and $h\in\mathcal{C}^{\infty}(X\times[0,1];\mathbb{R})$ a Hamiltonian on $X$. Then $L$ and $\phi^{1}(L)$ are Lagrangian cobordant, where $\phi^{t}$ is the time $t$-flow of the Hamiltonian vector field associated to $h$. This can be realised by the Hamiltonian suspension cobordism
$V:$ $\hskip 2.84526ptL\times[0,1]\longrightarrow X\times\mathbb{C}$
@@
which is constructed in detail in [bc1, Section 6.1]. More involved cobordisms can be constructed from general Lagrangian $k$-(anti)surgeries, see [haug_antisurgery].
If $dim(X)=2n$ and when one restricts to oriented Lagrangians, there is a cycle class map
$cyc:\hskip 5.69054pt\mathcal{C}ob^{or}(X)$ $\longrightarrow H_{n}(X;\mathbb{Z})$
@@
from which one can define the homologically trivial subgroup $\mathcal{C}ob^{or}(X;\mathbb{Z})_{0}$ appearing in the exact sequence
$\mathcal{C}ob(X)^{or}_{0}\hookrightarrow\mathcal{C}ob^{or}(X)\longrightarrow H_{n}(X;\mathbb{Z}).$
While proving the existence of a cobordism usually relies on exhibiting Lagrangian surgeries or Hamiltonian isotopies between Lagrangians, obstructions to their existence in the oriented case can be obtained by the presence of some symplectic flux bewteen homologically equivalent Lagrangians. This is the strategy employed by Arnold and Haug in [Arnold_Lag&Leg_cobordisms, haug_T2] to compute $\mathcal{C}ob(T^{*}S^{1})^{or}$ and $\mathcal{C}ob(T^{2})^{or}$, respectively.
@@
Let $L$ and $L^{\prime}$ be oriented Lagrangians in $M^{2n}$ with $L=L^{\prime}$ in $\mathcal{C}ob^{or}(X)$. Consider a smooth $(n+1)$-chain $\gamma$ in $M$ satisfying $\partial\gamma=L-L^{\prime}$. Then, for any closed $(n-1)$-form $\alpha$ on $M$,
@@
The equality $L=L^{\prime}$ in $\mathcal{C}ob(X)$ meants there exists Lagrangians $L_{1},\dots,L_{k}$ and a Lagrangian cobordism $V\subset M\times\mathbb{C}$ betweeen $(L_{1},\dots,L_{j},L,L_{j+1},\dots,L_{k})$ and $(L_{1},\dots,L_{i},L^{\prime},L_{i+1},\dots,L_{k})$. At the cost of making $V$ immersed, one can modify its cylindrical ends so that they terminate over a single fibre $\{z\}\in\mathbb{C}$. Denote this modified Lagrangian by $\tilde{V}$. Then $\tilde{V}\cup(\gamma\times\{z\})$ is an $(n+1)$-cycle in $M\times\mathbb{C}$. Let $\pi_{M}:M\times\mathbb{C}\longrightarrow M$ and $\pi_{\mathbb{C}}:M\times\mathbb{C}\longrightarrow\mathbb{C}$ be projections onto the first and second factor respectively. Notice that, if $r:\mathbb{C}\rightarrow\{z\}$ is a strong deformation retraction, then
@@
$=\int_{\gamma\times\{z\}}\pi_{M}^{*}(\alpha\wedge\omega)~$
@@
One can rephrase this Lemma by saying that, for any closed $(n-1)$-form $\alpha$ on $M$, there exists a flux-type morphism
(2.1) $\Theta_{\alpha}:$ $\hskip 8.53581pt\mathcal{C}ob^{or}(X)_{0}\longrightarrow\mathbb{R}/(\alpha\wedge\omega)(H_{n+1}(X))$
$\hskip 14.22636ptL\hskip 17.07164pt\mapsto\hskip 22.76219pt\int_{\gamma}\alpha\wedge\omega,$
@@
Let $T$ be an $n$-torus, and $\eta$ be a closed $1$-form on $T$. Let $T_{0}$ be the zero section, and $T_{\eta}$ the graph of $\eta$. Then $T_{0}=T_{\eta}$ in $\mathcal{C}ob^{or}(X)$ if and only if $\eta$ is exact.
@@
Notice that if $\eta$ is exact, the Hamiltonian suspension cobordism 2.2 between $T$ and $T_{\eta}$ is oriented. On the other hand, $\eta$ being non-exact implies the existence of a primitive class $\alpha=(z_{1},\dots,z_{n})\in H_{1}(T,\mathbb{Z})$ such that $\int_{\alpha}\eta\neq 0$. Such a class can be represented by a cycle $\gamma$ given by a rational line in $[0,1]^{n}/\sim$. A choice of supplementary plane to $\gamma$ in $[0,1]^{n}/\sim$ yields a diffeomorphism $T\cong S^{1}\times T^{n-1}$, which in turn gives a symplectomorphism
$\phi:(T^{*}T,\omega)\rightarrow(T^{*}(S^{1}\times T^{n-1}),\omega_{\gamma}\oplus\omega_{n-1}),$
with $\omega$, $\omega_{\gamma}$, and $\omega_{n-1}$ respectively denoting the standard symplectic forms on $T^{*}T$, $T^{*}S^{1}$, and $T^{*}T^{n-1}$.
@@
We wish to apply Lemma 2.4 by integrating over $A$ the $(n+1)$-form $\omega\wedge\phi^{*}vol\in\Omega^{n+1}(T^{*}T)$ over $A$, where $vol\in\Omega^{n-1}(T^{*}(S^{1}\times T^{n-1}))$ is the pullback by the natural projection map $\pi:T^{*}(S^{1}\times T^{n-1})\rightarrow T^{n-1}$ of an integral volume form on $T^{n-1}$. Notice that, because $H^{n+1}(T^{*}T^{n}\times\mathbb{C})$ is trivial, it is enough to show that this integral does not vanish.
$\int_{A}\omega\wedge\phi^{*}vol$ $=\int_{\phi_{*}A}(\omega_{\gamma}\oplus\omega_{n-1})\wedge vol$
$=\int_{\phi_{*}A}\omega_{\gamma}\wedge vol.$
Denoting by $\tilde{\pi}:\phi_{*}A\rightarrow T^{n-1}$ the restriction to $\phi_{*}A$ of the projection $\pi$, and by $y$ any point in $T^{n-1}$, we can apply fiberwise integration on each of the cylindrical fibres $\tilde{\pi}^{-1}(y)$ to obtain
$\int_{\phi_{*}A}\omega_{\gamma}\wedge vol$ $=\int_{T^{n-1}}\omega_{\gamma}(\tilde{\pi}^{-1}(y))vol.$
Notice now that $\omega_{\gamma}(\tilde{\pi}^{-1}(y))=\int_{\gamma}\eta$, therefore
$\int_{A}\omega\wedge\phi^{*}vol=vol(T^{n-1})\int_{\gamma}\eta\neq 0.$
@@
The result above does not hold for a general embedded Lagrangian torus $\iota:T^{n}\hookrightarrow X^{2n}$ in any symplectic manifold. By requiring that
$\iota^{*}:H^{1}(X;\mathbb{R})\longrightarrow H^{1}(T^{n};\mathbb{R})$
is onto, one can always find a non-zero integral over an $(n+1)$-chain bounding the Lagrangians $T^{n}$ and $T^{n}_{\eta}$ for some non-exact $\eta\in H^{1}(T^{n};\mathbb{R})$, as constructed in the proof of Lemma 2.6. However, this no longer translates to a non-zero flux because one might have $H_{n+1}(X;\mathbb{Z})\neq 0$. Counterexamples appear in recent work by Chassé-Leclercq [chassé2024weinsteinexactnessnearbylagrangians]. They construct Lagrangian tori in any symplectic manifold of dimension $\geq 6$ which admit arbitrarily Hausdorff-close disjoint elements in their Hamiltonian orbit. This follows from the characterisation of such orbits for product tori in $\mathbb{C}^{n}$ by Chekanov in [Chekanov1996LagrangianTI]. Crucially, these Lagrangian tori are not rational in the sense that $\omega(H_{2}(X,L))\neq t\mathbb{Z}$ for any $t\geq 0$. In dimension four, forthcoming results by Brendel–Kim announced by Brendel in a talk at ETH in October 2024 exhibit Hamiltonian isotopic Lagrangian tori $\iota:T\hookrightarrow(S^{2}\times S^{2},\omega\oplus t\omega)$ for $t\notin\mathbb{Q}$ and $\iota^{\prime}:T^{\prime}\hookrightarrow(S^{2}\times S^{2},\omega\oplus t\omega)$ such that $T^{\prime}\cong T_{\eta}$ for $\eta\neq 0\in H^{1}(T;\mathbb{R})$ in a Weinstein neighbourhood of $T$.
@@
In the interest of studying the Fukaya category of a symplectic manifold, one might want to consider Lagrangian branes. These are Lagrangians equipped with a grading, in order for their morphism spaces to be $\mathbb{Z}$-graded, and (twisted) Pin structures, which are used to orient the moduli spaces defining the $A_{\infty}$ operations, therefore are required unless one is content to work with $\mathbb{Z}_{2}$-coefficients.
The important implication here is that a graded Lagrangian automatically possesses a relative orientation with respect to some local system $\xi\subset\Lambda^{n}_{\mathbb{C}}TX$. Details can be found in [seidel_book, Section 11j] (recall that defining such gradings required the assumption $02c_{1}(X)=$ 0). When furthermore $c_{1}(X)=0$, in particular for the tori we are concerned with in this paper, one can choose local system to be trivial. Therefore these relative orientations are in fact honest orientations on the Lagrangians, and cobordisms of Lagrangian branes should preserve these orientations.
@@
Of course in the graded setting but with $c_{1}(X)$ $2$-torsion, one would still have a relative cycle class map
$cyc_{\xi}:\mathcal{C}ob(X)\longrightarrow$ $H_{n}(X;\xi)$
@@
and flux obstructions can be constructed analogously in this setting using the theory developped in [local_coefficients_steenrod], through a “local coefficient version” of 2.1.
@@ 3. Algebraic Lagrangian cobordisms
This Section uses some of the tropical background reviewed in 4.1 and 4.2 concerning tropical tori, as well as tropical curves and their Jacobians.
@@ 3.1. Motivation and definitions
As advertised in the introduction, our motivation for defining algebraic Lagrangian cobordisms is to provide an equivalence relation on Lagrangians mirror to algebraic equivalence of cycles. Because we use the language of Lagrangian correspondences to define and study algebraic Lagrangian cobordisms, the Lagrangians are only assumed to be immersed, not embedded. In fact, the Lagrangian Ceresa cycle 7 is immersed.
In this paper all of the algebraic varieties we consider are defined over $\mathbb{C}$. Let $Y$ be an algebraic variety, and $Z_{0}$, $Z_{\infty}$ two algebraic $k$-cycles in $Y$. The most frequent definition one will encounter in the literature is that $Z_{0}$ and $Z_{\infty}$ are said to be algebraically equivalent if they are two fibres of a flat family over a curve. However, an equivalent definition replaces the curve by any smooth projective variety; this is the version found in [fulton_intersection_theory, §10.3]. The definition which is most relevant to us replaces the curve by an abelian variety. Equivalence of these three definitions was proved by A. Weil in [weil_criteres_equivalence, Lemme 9] (translated in [abelian_varieties_lang, §III.1]), and the more general statement in modern algebro-geometric language is due to Achter–Casalaina-Martin–Vial in [parameter_spaces_alg_equiv].
@@
The cycles $Z_{0}$ and $Z_{\infty}$ are said to be algebraically equivalent if they are fibres over points $p,q\in B$ of a $(k+g)$-cycle $\Gamma$ in $Y\times B$, flat over $B$, where $B$ is an abelian variety of dimension $\dim(B)=g$ over $\mathbb{C}$.
@@
This will motivate our definition of algebraic Lagrangian cobordisms. We call symplectic polarised torus a symplectic manifold $X(B):=T^{*}B/T^{*}_{\mathbb{Z}}B$ where $B$ is a polarised tropical torus (tropical tori and polarisations thereof are introduced in Section 4).
Similarly as for Lagrangian cobordisms, one can choose a set of suitable Lagrangians on which to define the equivalence relation. Furthermore, if one endows these Lagrangians with additional decorations (eg orientation, local systems, etc), one should compatibly decorate the algebraic Lagrangian cobordism. Very generally, we will denote by $\mathcal{L}(X)$ a set of suitable (decorated) Lagrangians in $X$. As we have motivated at the end of the previous section, for our purposes this will mean closed oriented Lagrangians.
@@
Let $(L_{1},\dots,L_{k})$ and $(L_{1}^{\prime},\dots,L_{k^{\prime}}^{\prime})$ be two finite sets of Lagrangians in $\mathcal{L}(X)$. They are said to be algebraically Lagrangian cobordant if there exists a symplectic polarised torus $A:=T^{*}B/T^{*}_{\mathbb{Z}}B$ with points $p,q\in B$, and a Lagrangian correspondence $\Gamma\in\mathcal{L}(X\times A)$ such that
  • $\Gamma(F_{p})=L_{1}+\dots+L_{k}$ and $\Gamma(F_{q})=L^{\prime}_{1}+\dots+L^{\prime}_{k^{\prime}}$, where $F_{p}$ and $F_{q}$ are the fibres in $A$ at $p$ and $q$ respectively;
@@
    $\pi_{B}:X\times A\longrightarrow B$
    is the composition of the projection onto the second factor $X\times A\rightarrow A$ with the natural projection to the base $A\rightarrow B$.
Then we call $\Gamma$ an algebraic Lagrangian cobordism, or an algebraic Lagrangian correspondence between $(L_{1},\dots,L_{k})$ and $(L_{1}^{\prime},\dots,L_{k^{\prime}}^{\prime})$ and denote this $\Gamma:(L_{1},\dots,L_{k})\stackrel{{\scriptstyle A_{p,q}}}{{\rightsquigarrow}}(L_{1}^{\prime},\dots,L_{k^{\prime}}^{\prime})$ or simply $\Gamma:(L_{1},\dots,L_{k})\stackrel{{\scriptstyle A}}{{\rightsquigarrow}}(L_{1}^{\prime},\dots,L_{k^{\prime}}^{\prime})$. From this we define the algebraic Lagrangian cobordism group of $X$ as
$algCob(X):=\oplus_{L\in\mathcal{L}}\mathbb{Z}\cdot L/\sim,$
@@
We will focus our attention on the oriented Lagrangian cobordism group, which we denote $algCob^{or}(X)$. Here the Lagrangians considered are oriented, and any algebraic Lagrangian cobordism is also oriented with equalities of the type $\Gamma(F)=L$ preserving the orientation (notice this requires a choice of orientation for fibres $F$ of $X(B)$.
@@
$algCob^{or}(X)\longrightarrow H_{n}(X;\mathbb{Z}).$

### v2

@@ 2.2.4. Tropical flux
The notion of Lagrangian cobordism was originally introduced by Arnold in [Arnold_Lag_Leg_cobordisms].
@@
Two finite sets of Lagrangians $(L_{1},\dots,L_{k^{+}})$ and $(L^{\prime}_{1},\dots,L^{\prime}_{k^{-}})$ in $M$ are Lagrangian cobordant if there exists a Lagrangian $V$ in $M\times\mathbb{C}$ and a compact subset $K\subset M\times\mathbb{C}$ such that
@@
for pairwise distinct real numbers $\{x_{i}\}$ and $\{x^{\prime}_{j}\}$. The Lagrangian cobordism group of $M$, $\mathcal{C}ob(M)$, is then
$\mathcal{C}ob(M):=\bigoplus_{L\in M}\mathbb{Z}\cdot L/\sim,$
@@
This definition can be modified to fit the desired set-up by adding assumptions on the finite sets of Lagrangians and the cobordisms themselves with a compatibility condition on the ends; in this way one can upgrade Lagrangians to Lagrangian branes, require exactness, monotonicity, etc. Most relevant to our considerations is the oriented Lagrangian cobordism group, which we denote by $\mathcal{C}ob^{or}(M)$.
A slight variation of Lagrangian cobordisms are cylindrical Lagrangian cobordisms, introduced by Sheridan–Smith in [nick_ivan1]. These differ from Definition 3.1 in that $V$ is a Lagrangian in $M\times\mathbb{C}^{*}$, which outside of a compact subset in $\mathbb{C}^{*}$ carries trivially fibered finite sets of Lagrangians in $M$ over radial rays. This variation appears naturally from mirror symmetry, where cylindrical cobordisms provide the more natural mirror to rational equivalence.
@@
Let $L\subset M$ be a Lagrangian, and $h\in\mathcal{C}^{\infty}(M\times[0,1];\mathbb{R})$ a Hamiltonian on $M$. Then $L$ and $\phi^{1}(L)$ are Lagrangian cobordant, where $\phi^{t}$ is the time $t$-flow of the Hamiltonian vector field associated to $h$. This can be realised by the Hamiltonian suspension cobordism
$V:$ $\hskip 2.84526ptL\times[0,1]\longrightarrow M\times\mathbb{C}$
@@
which is constructed in detail in [bc1, Section 6.1]. More involved cobordisms can be constructed from general Lagrangian kk-(anti)surgeries, see [haug_antisurgery].
If $dim(M)=2n$ and when one restricts to oriented Lagrangians, there is a cycle class map
$cyc:\hskip 5.69054pt\mathcal{C}ob^{or}(M)\longrightarrow$ $H_{n}(M;\mathbb{Z})$
@@
from which one can define the homologically trivial subgroup $\mathcal{C}ob^{or}(M;\mathbb{Z})_{0}$ appearing in the exact sequence
$\mathcal{C}ob(M)^{or}_{0}\hookrightarrow\mathcal{C}ob^{or}(M)\longrightarrow H_{n}(M;\mathbb{Z}).$
While proving the existence of a cobordism usually relies on exhibiting Lagrangian surgeries or Hamiltonian isotopies between Lagrangians, obstructions to their existence in the oriented case can be obtained by the presence of some symplectic flux between homologically equivalent Lagrangians. This is the strategy employed by Arnold and Haug in [Arnold_Lag_Leg_cobordisms, haug_T2] to compute $\mathcal{C}ob(T^{*}S^{1})^{or}$ and $\mathcal{C}ob(T^{2})^{or}$, respectively.
@@
Let $L$ and $L^{\prime}$ be oriented Lagrangians in $M^{2n}$ with $L=L^{\prime}$ in $\mathcal{C}ob^{or}(M)$. Consider a smooth $(n+1)$-chain $\gamma$ in $M$ satisfying $\partial\gamma=L-L^{\prime}$. Then, for any closed (n1)(n-1)-form $\alpha$ on $M$,
@@
The equality $L=L^{\prime}$ in $\mathcal{C}ob(M)$ means there exists Lagrangians $L_{1},\dots,L_{k}$ and a Lagrangian cobordism $V\subset M\times\mathbb{C}$ betweeen $(L_{1},\dots,L_{j},L,L_{j+1},\dots,L_{k})$ and $(L_{1},\dots,L_{i},L^{\prime},L_{i+1},\dots,L_{k})$. At the cost of making $V$ immersed, one can modify its cylindrical ends so that they terminate over a single fibre $\{z\}\in\mathbb{C}$. Denote this modified Lagrangian by $\tilde{V}$. Then $\tilde{V}\cup(\gamma\times\{z\})$ is an $(n+1)$-cycle in $M\times\mathbb{C}$. Let $\pi_{M}:M\times\mathbb{C}\longrightarrow M$ and $\pi_{\mathbb{C}}:M\times\mathbb{C}\longrightarrow\mathbb{C}$ be projections onto the first and second factor respectively. Notice that, if $r:\mathbb{C}\rightarrow\{z\}$ is a strong deformation retraction, then
@@
$=\int_{\gamma\times\{z\}}\pi_{M}^{*}(\alpha\wedge\omega)$
@@
One can rephrase this Lemma by saying that, for any closed (n1)(n-1)-form $\alpha$ on $M$, there exists a flux-type morphism
(3.1) $\Theta_{\alpha}:$ $\hskip 8.53581pt\mathcal{C}ob^{or}(M)_{0}\longrightarrow\mathbb{R}/(\alpha\wedge\omega)(H_{n+1}(M;\mathbb{Z}))$
$\hskip 28.45274ptL\hskip 22.76219pt\mapsto\hskip 28.45274pt\int_{\gamma}\alpha\wedge\omega,$
@@
Let $T$ be an $n$-torus, and $\eta$ be a closed $1$-form on $T$. Let $T_{0}$ be the zero section, and $T_{\eta}$ the graph of $\eta$. Then $T_{0}=T_{\eta}$ in $\mathcal{C}ob^{or}(M)$ if and only if $\eta$ is exact.
@@
Notice that if $\eta$ is exact, the Hamiltonian suspension cobordism from Example 3.2 between $T$ and $T_{\eta}$ is oriented. On the other hand, $\eta$ being non-exact implies the existence of a primitive class $z=(z_{1},\dots,z_{n})\in H_{1}(T;\mathbb{Z})$ such that $\int_{z}\eta\neq 0$. Such a class can be represented by a cycle $Z$ given by a rational line in $0,1]^{n}/\sim$. A choice of supplementary plane to $Z$ in $0,1]^{n}/\sim$ yields a diffeomorphism $T\cong S^{1}\times T^{n-1}$, which in turn gives a symplectomorphism
$\phi:(T^{*}T,\omega)\rightarrow(T^{*}(S^{1}\times T^{n-1}),\omega_{Z}\oplus\omega_{n-1}),$
with $\omega$, $\omega_{Z}$, and $\omega_{n-1}$ respectively denoting the standard symplectic forms on $T^{*}T$, $T^{*}S^{1}$, and $T^{*}T^{n-1}$.
@@
We wish to apply Lemma 3.4 by integrating over $\gamma$ the $(n+1)$-form $\omega\wedge\phi^{*}vol\in\Omega^{n+1}(T^{*}T)$, where $vol\in\Omega^{n-1}(T^{*}(S^{1}\times T^{n-1}))$ is the pullback by the natural projection map $\pi:T^{*}(S^{1}\times T^{n-1})\rightarrow T^{n-1}$ of an integral volume form on $T^{n-1}$. Notice that, because $H^{n+1}(T^{*}T^{n}\times\mathbb{C})$ is trivial, it is enough to show that this integral does not vanish.
$\int_{\gamma}\omega\wedge\phi^{*}vol$ $=\int_{\phi_{*}\gamma}(\omega_{Z}\oplus\omega_{n-1})\wedge vol$
$=\int_{\phi_{*}\gamma}\omega_{Z}\wedge vol.$
This last integral can be simplified as follows. Let $\tilde{\pi}:\phi_{*}\gamma\rightarrow T^{n-1}$ be the restriction to $\phi_{*}\gamma$ of the projection $\pi$ (which is surjective), and $y$ any point in $T^{n-1}$. Each fibre
$\tilde{\pi}^{-1}(y)=\{(s,y,t(\phi^{-1})^{*}\eta(s,y)):s\in Z,t\in[0,1]\}$
is a compact cylinder parametrised by $s\in Z\simeq S^{1}$ and $t\in[0,1]$. Applying fibrewise integration we get
$\int_{\phi_{*}\gamma}\omega_{Z}\wedge vol$ $=\int_{T^{n-1}}\left(\int_{\tilde{\pi}^{-1}(y)}\omega_{Z}\right)vol.$
Notice now that $\int_{\tilde{\pi}^{-1}(y)}\omega_{Z}=\int_{Z}\eta$ is independent of $y\in T^{n-1}$, therefore
$\int_{\gamma}\omega\wedge\phi^{*}vol=vol(T^{n-1})\int_{Z}\eta\neq 0.$
@@
The result above does not hold for a general embedded Lagrangian torus $\iota:T^{n}\hookrightarrow M^{2n}$ in any symplectic manifold. By requiring that
$\iota^{*}:H^{1}(M;\mathbb{R})\longrightarrow H^{1}(T^{n};\mathbb{R})$
is onto, one can always find a non-zero integral over an $(n+1)$-chain bounding the Lagrangians $T^{n}$ and $T^{n}_{\eta}$ for some non-exact $\eta\in H^{1}(T^{n};\mathbb{R})$, as constructed in the proof of Lemma 3.6. However, this no longer translates to a non-zero flux because one might have $H_{n+1}(M;\mathbb{Z})\neq 0$. Counterexamples appear in recent work by Chassé-Leclercq [chassé2024weinsteinexactnessnearbylagrangians]. They construct Lagrangian tori in any symplectic manifold of dimension $\geq 6$ which admit arbitrarily Hausdorff-close disjoint elements in their Hamiltonian orbit. This follows from the characterisation of such orbits for product tori in $\mathbb{C}^{n}$ by Chekanov in [Chekanov1996LagrangianTI]. Crucially, these Lagrangian tori are not rational in the sense that $\omega(H_{2}(M,L))\neq t\mathbb{Z}$ for any $t\geq 0$. In dimension four, Brendel–Kim exhibit Hamiltonian isotopic Lagrangian tori $\iota:T\hookrightarrow(S^{2}\times S^{2},\omega\oplus t\omega)$ for $t\in\mathbb{R}$ and $\iota^{\prime}:T^{\prime}\hookrightarrow(S^{2}\times S^{2},\omega\oplus t\omega)$ such that $T^{\prime}\cong T_{\eta}$ for $\eta\neq 0\in H^{1}(T;\mathbb{R})$ in a Weinstein neighbourhood of $T$ [brendel_kim].
@@
In the interest of studying the Fukaya category of a symplectic manifold, one might want to consider Lagrangian branes. These are Lagrangians equipped with a grading in order for their morphism spaces to be $\mathbb{Z}$-graded, as well as (twisted) Pin structures which are used to orient the moduli spaces defining the $A_{\infty}$ operations, and are therefore required unless one is content to work with $\mathbb{Z}_{2}$-coefficients.
The important implication here is that a graded Lagrangian automatically possesses a relative orientation with respect to some local system $\xi\subset\Lambda^{n}_{\mathbb{C}}TM$. Details can be found in [seidel_book, Section 11j] (recall that defining such gradings required the assumption $2c_{1}(M)=0$). When furthermore $c_{1}(M)=0$, in particular for the tori we are concerned with in this paper, one can choose this local system to be trivial. Therefore these relative orientations are in fact honest orientations on the Lagrangians, and cobordisms of Lagrangian branes should preserve these orientations.
@@
Of course in the graded setting but with $c_{1}(M)$ 22-torsion, one would still have a relative cycle class map
$cyc_{\xi}:\mathcal{C}ob(M)\longrightarrow$ $H_{n}(M;\xi)$
@@
and flux obstructions could be constructed analogously in this setting using the theory developped in [local_coefficients_steenrod], through a “local coefficient version” of the morphism in Equation (3.1).
@@ 4. Algebraic Lagrangian cobordisms
This Section uses some of the tropical background reviewed in Sections 2.1.1 and 2.1.2 concerning tropical tori and tropical curves.
@@ 4.1. Motivation and definitions
As advertised in the introduction, our motivation for defining algebraic Lagrangian cobordisms is to provide an equivalence relation on Lagrangians mirror to algebraic equivalence of cycles. Because we use the language of Lagrangian correspondences to define and study algebraic Lagrangian cobordisms, we work with immersed Lagrangians.
Let $Y$ be an algebraic variety over $\mathbb{C}$, and $_{0}, ZZ_{\infty}$ two algebraic $k$-cycles in $Y$. The most frequent definition found in the literature is that $Z_{0}$ and $Z_{\infty}$ are said to be algebraically equivalent if they are two fibres of a flat family over a curve. However, an equivalent definition replaces the curve by any smooth projective variety; this is the version found in [fulton_intersection_theory, §10.3]. The definition which is most relevant to us replaces the curve by an abelian variety. Equivalence of these three definitions was proved by A. Weil in [weil_criteres_equivalence, Lemme 9] (translated in [abelian_varieties_lang, §III.1]), and in modern algebro-geometric language by Achter–Casalaina-Martin–Vial in [parameter_spaces_alg_equiv].
@@
The cycles $Z_{0}$ and $Z_{\infty}$ are said to be algebraically equivalent if they are fibres over points $p,q\in B$ of a $(k+g)$-cycle $\Gamma$ in $Y\times B$, flat over $B$, where $B$ is an abelian variety of dimension $g$ over $\mathbb{C}$.
@@
We call symplectic polarised torus a symplectic manifold $X(B):=T^{*}B/T^{*}_{\mathbb{Z}}B$ where $B$ is a polarised tropical torus (tropical tori and polarisations thereof are introduced in Section 2.1). Recall that a Lagrangian correspondence $\Gamma$ between symplectic manifolds $M_{1}$ and $M_{2}$ is a Lagrangian in $\overline{M_{1}}\times M_{2}$. Given a Lagrangian $L_{1}$ in $M_{1}$, the geometric composition $\Gamma(L_{1})$ of $\Gamma$ with $L_{1}$ is $\pi_{M_{2}}(\Gamma\cap(L_{1}\times M_{2}))\subset M_{2}$. Possibly after perturbing $\Gamma$ or $L_{1}$ by a Hamiltonian isotopy, the intersection can be made transverse, and in this case $\Gamma(L_{1})$ is a immersed Lagrangian (see e.g. [ww_quilted_floer_cohomology][Lemma 2.0.5]).
Similarly as for Lagrangian cobordisms, one can choose a set of suitable Lagrangians on which to define the equivalence relation. Furthermore, if one endows these Lagrangians with additional decorations (e.g. orientation, local systems, etc), one should compatibly decorate the algebraic Lagrangian cobordism. Very generally, we will denote by $\mathcal{L}(M)$ a set of suitable (decorated) Lagrangians in the symplectic manifold $M$. As we have motivated at the end of the previous section, for our purposes this will mean closed oriented Lagrangians.
@@
Let $(L_{1},\dots,L_{k})$ and $(L_{1}^{\prime},\dots,L_{k^{\prime}}^{\prime})$ be two finite sets of Lagrangians in $\mathcal{L}(M)$. They are said to be algebraically Lagrangian cobordant if there exists a symplectic polarised torus $X(B):=T^{*}B/T^{*}_{\mathbb{Z}}B$ with points $p,q\in B$, and a Lagrangian correspondence $\Gamma\in\mathcal{L}(\overline{X(B)}\times M)$ such that
  • $\Gamma(F_{p})=L_{1}+\dots+L_{k}$ and $\Gamma(F_{q})=L^{\prime}_{1}+\dots+L^{\prime}_{k^{\prime}}$, where $F_{p}$ and $F_{q}$ are the fibres in $X(B)$ at $p$ and $q$ respectively;
@@
    $\pi_{B}:M\times X(B)\longrightarrow B$
    is the composition of the projection onto the second factor $M\times X(B)\rightarrow X(B)$ with the natural projection to the base $X(B)\rightarrow B$.
Then we call $\Gamma$ an algebraic Lagrangian cobordism, or an algebraic Lagrangian correspondence between $(L_{1},\dots,L_{k})$ and $(L_{1}^{\prime},\dots,L_{k^{\prime}}^{\prime})$ and denote this $\Gamma:(L_{1},\dots,L_{k})\stackrel{{\scriptstyle X(B)_{p,q}}}{{\rightsquigarrow}}(L_{1}^{\prime},\dots,L_{k^{\prime}}^{\prime})$ or simply $\Gamma:(L_{1},\dots,L_{k})\stackrel{{\scriptstyle X(B)}}{{\rightsquigarrow}}(L_{1}^{\prime},\dots,L_{k^{\prime}}^{\prime})$. From this we define the algebraic Lagrangian cobordism group of $M$ as
$algCob(M):=\oplus_{L\in\mathcal{L}}\mathbb{Z}\cdot L/\sim,$
@@
We will focus our attention on the oriented Lagrangian cobordism group, which we denote $algCob^{or}(M)$. Here the Lagrangians considered are oriented, and any algebraic Lagrangian cobordism is also oriented with equalities of the type $\Gamma(F)=L$ respecting the orientation (notice this requires a choice of orientation for fibres $F$ of $X(B)$).
@@
$algCob^{or}(M)\longrightarrow H_{n}(M;\mathbb{Z}).$